OpenStudy (anonymous):

Find the value of $10000 at the end of one year if it is invested in an account that has an interest rate of 4.95% and is compounded in accordance with the rules below. (a). compounded monthly (b). compounded daily (assuming a 365- day year) (c). compounded quarterly

OpenStudy (phoenixfire):

Do you know the compound interest formula?

OpenStudy (phoenixfire):

\[F=P(1+{r \over n})^{nt}\]F=Future amount P=Initial amount r=annual interest rate n=number of times compounded in a year t=number of years. So, for (a), compounded monthly. n=12, t=1, P=10000, r=0.0495. Substitute all that in and you'll have your F, the future amount.

OpenStudy (anonymous):

Actually they are giving me the formula A=P(1+r/m)mt

OpenStudy (anonymous):

\[10000(1+0.0495\div12)12\]

OpenStudy (phoenixfire):

Same thing, different letters.

OpenStudy (anonymous):

ok. SO HOW CAN I FIND THE ANSWER BECAUSE WHAT IM GETTING DOESNT SEEM CORRECT TO WHAT THE EXAMPLE IS SHOWING ME

OpenStudy (anonymous):

I pretty much have found the computing formula to find each steps but unable to get verifiable answer

OpenStudy (anonymous):

because in the examples they use 15000 and the answer turns out to be $15696.91

OpenStudy (phoenixfire):

\[A=P(1+{r \over m})^{mt}\]\[A=10000(1+{0.0495\over 12})^{12}\]\[A=10000(1.05)\]\[A=10506.4\] Obviously I've done rounding where I shouldn't have, but that's the procedure.

OpenStudy (anonymous):

how did they get that answers from using the formula because i know for m- i would put 12

OpenStudy (phoenixfire):

What are the values they use in the examples?

OpenStudy (anonymous):

they used 15000-p, 0.0455- r, and 12- m

OpenStudy (phoenixfire):

Right so, put those values in in the formula. \[A=15000(1+{0.0455\over 12})^{12}\]

OpenStudy (anonymous):

yes

OpenStudy (phoenixfire):

And you get: 15696.91

OpenStudy (anonymous):

yes thats what they got

OpenStudy (anonymous):

how did u reach to that?

OpenStudy (phoenixfire):

By working out the above equation. Lets do it step by step... Upcoming Ugly numbers!!! 0.0455/12=0.00379166666666666666666666666667 That+1=1.00379166666666666666666666666667 that ^ 12=1.0464609601126369220378253015612 that * 15000=15696.914401689553830567379523418 with some rounding, you get 15696.91.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

let me try it myself and i will let you know

OpenStudy (phoenixfire):

Order of operations. Brackets Exponents Division Multiplication Addition Subtraction or easily remembering: BEDMAS. Also, remember ab^k is NOT equal (ab)^k. the difference is that ab^k is a*(b^k) whereas (ab)^k is (a*b) ^ k... there's a different order that you apply the maths in; which result in a different value at the end.

OpenStudy (anonymous):

ok

OpenStudy (phoenixfire):

Good luck.

OpenStudy (anonymous):

it didnt work the way u did it when i plugged everything myself into the equation It give me 10,506.39 and i got 2690666.1

OpenStudy (phoenixfire):

Work it out in steps. 0.0495/12=a 1+a=b b^12=c P*c=F